OpenStudy (anonymous):

If s=siny, y ∈ [-pi/2,pi/2], what is dy/ds?

OpenStudy (anonymous):

I don't even understand what it's asking... So it wants to know the derivative of y with respect to the derivative of s? So s`=cosy? cosy/s`=dy/ds??? Completely guessing, honestly have no idea how to proceed with this...

OpenStudy (anonymous):

\[dy/ds=1/\sqrt{1-s ^{2}}\]

myininaya (myininaya):

s=sin(y) => y=arcsin(s) but the inverse sine only exists if we restrict the domain of sine from [-pi/2,pi/2] so let's look at this s=sin(y) we want to find the derivative of y with respect to s so (s)'=1 and (y)'=dy/ds => (sin(y))'=dy/ds * cos(y) so we have 1=dy/ds * cos(y) solving for dy/ds we have 1/cos(y)=dy/ds dy/ds=1/cos(y) assume we have a right triangle Recall sin(y)=s=s/1 (=opp/hyp) now we need to find the adj side we can use the pythagorean thm to do this adj=sqrt(1-s^2) so we can rewrite dy/ds in terms of s dy/ds=1/sqrt(1-s^2) but remember y =arcsin(s) so we actually found the derivative of arcsin(s) with the restriction that -pi/2<=arcsin(s)<=pi/2